# An intense and controlled beam of S() atoms

### v1

@@ I Introduction
In this work, we demonstrate the production of a high-intensity, velocity-controlled beam of electronically excited S() atoms using a 3-m-long multistage Zeeman decelerator. The sulfur atoms are generated via photolysis of CS2 at 199.65 nm, which efficiently produces both S() and ground-state S() atoms. We show that both can be manipulated using the decelerator, and that temporal separation between these states can be achieved by decelerating the S atoms. The resulting S() beam has a well-defined velocity, narrow velocity spreads, and a high quantum-state purity if needed, making it ideally suited for high-resolution scattering experiments. As a proof of concept, we present an elastic scattering experiment involving collisions between S() and argon (Ar) atoms. The observed scattering signals confirm that the beam density is sufficient for detailed collision experiments.
@@ III.2 Quantum-state distributions
In hybrid mode, the S() and S() packets exit the decelerator simultaneously, and no separation of atoms in both quantum states is possible. However, in deceleration mode, the S() packet arrives later than the S() packet due to its larger magnetic moment and therefore stronger interaction with the magnetic fields. By selecting the arrival time corresponding to the peak of the S() packet, the quantum-state purity of the packet can be significantly improved. Note that, in a similar way, a high-purity S() packet could be produced as well.
@@ IV Conclusions
We have demonstrated the successful production of a high-intensity and velocity-controlled beam of electronically excited S() atoms using our multistage Zeeman decelerator. Following photolysis of CS2, both S() and ground-state S() atoms are formed, and both can be manipulated using the decelerator. By operating the decelerator in deceleration mode, we achieved temporal separation between S() and S() atoms, enabling the generation of a S-atom beam with high quantum-state purity when required.
Furthermore, we showed that the density of the S() beam is sufficient to perform elastic scattering experiments, as evidenced by our VMI image obtained for S() + Ar collisions. This is an excellent starting point for investigating reactive scattering between S() and H2 (including its isotopologues), as well as electronic quenching of S() in collisions with various partners, although it should be noted that the cross sections for these processes are generally expected to be smaller than those for the elastic S() + Ar collisions investigated here. Our experimental setup allows for precise control over the collision energy, making it well suited for exploring these processes under well-defined conditions.

### v2

# Intense and controlled beam of S() atoms

@@ I Introduction
In this work, we demonstrate the production of a high-intensity, velocity-controlled beam of electronically excited S() atoms using a 3-m-long multistage Zeeman decelerator. The sulfur atoms are generated via photolysis of CS2 at 199.65 nm, which efficiently produces both S() and ground-state S() atoms. We show that both can be manipulated using the decelerator, and that temporal separation between these states can be achieved by decelerating the S atoms. The resulting S() beam has a well-defined velocity, narrow velocity spreads, and an enhanced quantum-state purity if needed, making it ideally suited for high-resolution scattering experiments. As a proof of concept, we present an elastic scattering experiment involving collisions between S() and argon (Ar) atoms. The observed scattering signals confirm that the beam density is sufficient for detailed collision experiments.
@@ III.2 Quantum-state distributions
In hybrid mode, the S() and S() packets exit the decelerator simultaneously, and it is not possible to separate the atoms in the two quantum states. However, in deceleration mode, the S() packet arrives later than the S() packet due to its larger magnetic moment and therefore stronger interaction with the magnetic fields. By selecting the arrival time corresponding to the peak of the S() packet, the quantum-state purity of the packet can be significantly improved. The purity could be further enhanced by employing a heavier seed gas, which lowers the beam velocity and increases the effectiveness of deceleration, thereby improving the temporal separation between the different states. This approach, however, would also reduce the overall S() density, as discussed in section III.1. Note that, in a similar way, a high-purity S() packet could be produced as well.
@@ IV Conclusions
We have demonstrated the successful production of a high-intensity and velocity-controlled beam of electronically excited S() atoms using our multistage Zeeman decelerator. Following photolysis of CS2, both S() and ground-state S() atoms are formed, and both can be manipulated using the decelerator. By operating the decelerator in deceleration mode, we achieved temporal separation between S() and S() atoms, enabling the generation of a S-atom beam with enhanced quantum-state purity when required.
Furthermore, we showed that the density of the S() beam is sufficient to perform elastic scattering experiments, as evidenced by our VMI image obtained for S() + Ar collisions. This is an excellent starting point for investigating reactive scattering between S() and H2 (including its isotopologues), as well as electronic quenching of S() in collisions with various partners. Since cross sections for such processes could be one to two orders of magnitude smaller than those for elastic scattering, their measurement will be considerably more challenging. Nevertheless, the precise control over collision energy provided by our setup would make it well suited to explore these processes under well-defined conditions.
@@ V Acknowledgements
VI Data availability
The data that support the findings of this article are openly available [52].